Fone Zone Corporation manufactures telephones. Recently, the company produced a batch of 570 defective telephones at a cost of $8,700. Fone Zone can sell these telephones as scrap for $10 each. It can also rework the entire batch at a cost of $6,200, after which the telephones could be sold for $21 per unit.

If Fone Zone reworks the defective telephones, by how much will its operating income change?

Multiple Choice

  • Increase by $5,770

  • Increase by $6,270

  • Decrease by $2,930

  • Decrease by $5,770
Answer #1
Operating income changes
increase in operating income = sale -cost
cost = 6200
sale = 570*21 = 11970
Increase in operating income = 11970-6200
increase in operating income = 5770
Option " increase by $5770" is Correct

Please give a thumbs up if it is helpful & let me know if any doubt